Noah’s Ark  Scale 1:1000  Source: Genesis 6,14-16

↑ inscription onto a side wall

← window of transparent film

to be folded from the table plane to the upper direction to be folded from the table plane to the lower direction Handling: Norbert Südland, Aalen (Germany)

[figure] [figure] [figure] [figure] [figure] [figure] [figure] [figure] [figure] [figure] [figure] [figure] [figure] [figure] [figure] [figure]